# Nav-SCOPE: Swarm Robot Cooperative Perception and Coordinated Navigation

### v2

@@ Abstract
This paper proposes a lightweight systematic solution for multi-robot coordinated navigation with decentralized cooperative perception. An information flow is first created to facilitate real-time observation sharing over unreliable ad-hoc networks. Then, the environmental uncertainties of each robot are reduced by interaction fields that deliver complementary information. Finally, path optimization is achieved, enabling self-organized coordination with effective convergence, divergence, and collision avoidance. Our method is fully interpretable and ready for deployment without gaps. Comprehensive simulations and real-world experiments demonstrate reduced path redundancy, robust performance across various tasks, and minimal demands on computation and communication.
@@ I Introduction
Robot swarms have shown great potential in recent studies. From centralized control fleets [1, 2] to self-organized teams [3, 4], robots can now operate automatically in the real world. However, despite the significant progress, most solutions only use each robot’s local perception independently for navigation, while few works have ever utilized real-time cooperative perception of the swarm to optimize the planning process. Consequently, robots’ coordination is often confined to leader-follower [5, 6] or collision avoidance [4, 7] due to limited information flow within swarm systems, in which robots often encounter unexpected obstructions or unnecessary detours.
The advantages of cooperative perception are obvious: a wider field of view, more details of the unknown environment, and that lead to better paths with observation fusion. The challenges include but are not limited to informative environmental representation, decision-making with fused observation, on-board processing latency, and minimal unreliable communication. The first two are method requirements, and the latter two are system constraints.
@@
This paper proposes Nav-SCOPE: Swarm COoperative PErception and coordinated Navigation. This is an extremely lightweight systematic solution that addresses the above contradiction. In this system, each robot is used as a telescope of neighboring robots for an extended field of view. When robots decide their future paths, they will use the fused observation to choose better directions with fewer detours or conflicts. Experimental results show self-organized paths and better swarm automation. To the best of our knowledge, Nav-SCOPE is the first interpretable navigation solution with real-time cooperative perception, and it does not require a mapping process or external localization. The paper makes the following contributions:
@@
  • The navigation process is detailed for better coordination and less redundancy in the path. Parameters are determined theoretically by robot physical properties or extremely small samples.
  • The proposed Nav-SCOPE is deployed in fully autonomous swarms. Simulations and real-world experiments are performed in challenging environments.
@@ II Related Works
Numerous studies have focused on local planning since the advent of mobile robots. In the early stages, autonomous navigation was first realized in individual robots. Reactive strategies, such as Potential Fields (PFs) [8] and Bugs [9], are direct and lightweight solutions. Planners based on voxel maps [10, 11] can provide better local trajectories with more environmental details. These two methods do not require a prebuilt map, as a Simultaneous Localization and Mapping (SLAM) [12] process before planning execution is not suitable for many tasks and cases.
@@
As robot systems grow to fleets and swarms, critical issues of coordination between robots need to be solved. Centralized control is a typical way to regulate robot paths. For example, [1, 2] used central computing stations to achieve flock flight on multiple quadrotors. However, these approaches require ideal communication and external localization, making them vulnerable in scaled-up scenarios [13]. [3] used swarm gradient bug algorithm (SGBA) to achieve distributed control, but still needed an external beacon to provide a priori environmental information, which limits the scalability of the algorithm.
@@
Decentralized systems can be more versatile in large unknown environments. The more information utilized, the better performance a system would have. Currently, it is common to use mutual positioning and trajectory sharing to achieve self-organization and swarm navigation. Consequently, recent studies have provided two types of solutions. Leader-follower structures achieve synergic formation towards the same goal [14, 5, 6], while collision avoidance structures optimize robot paths to reduce negative interactions in various tasks [3, 15, 4, 7]. A typical state-of-the-art (SOTA) solution is EGO-Swarm-v2 [4], which performed joint optimizations of spatial-temporal trajectories and was tested in wild bamboo forests.
@@
Cooperative perception offers external environmental information to each robot, thereby holding significant potential to enhance swarm navigation and improve performance. However, these information flows are not utilized in current solutions. [16] proposed a Graph Neural Network (GNN) framework for collaborative perception, but its navigation process only performed well in grid maps [17], suffering from data collection and sim-to-real gaps. Nav-SCOPE is designed to solve these challenges and be deployed in real-world applications.
III Formulation of Nav-SCOPE
The system architecture of Nav-SCOPE is illustrated in Fig. 1. In this section, we will first introduce the information flow among swarm robots using cooperative perception. The coordinated navigation facilitated by observation fusion is then detailed.
III-A Information Flow from Cooperative Perception
The information flow within the swarm is formed by distributed sharing of robot environmental observation, including informative encoding from transmitter, and reconstruction at receiver. Consider a robot $R_{a}$ with a field of view (FOV) $\Theta$ and a width $r_{0}$. Its observation is defined as a depth sequence $s_{a}(n)$ with $M$ uniform samples within the FOV. To avoid $R_{a}$ from obstacles, we use a safety sector to evaluate the occupation of a desired direction, as shown in Fig. 2(a). The sector splits the FOV into four quadrants: navigation scope ($Q_{1}$, $Q_{2}$), and peripheral vision ($Q_{3}$, $Q_{4}$). The sector angle $\alpha$ and planning distance $l_{th}$ are given by:
@@ III-A Information Flow from Cooperative Perception
where $r$ is the safety margin from nearby obstacles. The navigation scope scans to evaluate all the directions within $s_{a}(n)$ before transmission. This process can be swiftly performed by low-pass filtering based on the Fast Fourier Transform (FFT), as demonstrated in a previous study [18]. The cutoff frequency of the filter can be decided directly by robot properties above:
$f_{c}=\frac{\Theta}{M\alpha}.$ (2)
The filtered observation $y_{a}(n)$ is further encoded into byte-level features $\mathbf{\Sigma}_{a}$ through the calculation of local extrema:
$\mathbf{\Sigma}_{a}=\left\{(\phi_{a,j},d_{a,j})|\Delta y_{a}(j)\Delta y_{a}(j+1)<0\right\},$ (3)
where $\phi_{a,j}$ and $d_{a,j}$ represent the directions and distances of the extrema from $R_{a}$. The directions of local maxima indicate possible open space, while the directions of local minima indicate potential obstacles. In this way, the computation cost is significantly reduced by FFT into log-linear level, and the communication load is minimal for unreliable networks. At the receiver, the observations are approximately reconstructed as polygons by connecting extrema vertices. Consequently, the information flow within the swarm is established. To strengthen the information flow, we made further improvements in three aspects:
@@
Thus, communication data are minimized while essential information is maintained. Reliability is guaranteed when using unreliable ad-hoc networks.
Figure 2: (a) Protective model and observation quadrants. In this case, the robot can keep a safety margin $r$ from the obstacle $B$. (b) Sampling process to determine the optimal $w_{i}$. (c) Magnitude curves of the interaction fields. (d) Performance comparison on simulation with different $w_{i}$.
III-B Coordinated Navigation and Path Optimization
To facilitate seamless coordination, we associate the information flow with an influence field to achieve probabilistic fusion. Here, the neighboring robots perform as telescopes to reduce environmental uncertainties, in other words, help to find open spaces with fewer obstacles towards the target. We use forces to indicate the level of impact. Individual intelligence is also formed for the fusion decisions.
III-B1 Interaction fields
They are connections among robots to achieve coordination and path optimization. Consider a case where $R_{a}$ observes its neighbor $R_{i}$. The influence of $R_{i}$ is closely related to its observation features $\Sigma_{i}$ and their relative distances $\|\vec{r}_{ai}\|$. The potential $V_{ai}$ of the interaction combines the effect of coordination convergence $V^{att}_{ai}$ and collision avoidance $V^{rep}_{ai}$:
@@ III-B1 Interaction fields
where $\sigma_{a}$ is the reference distance, marking the transition from repulsion to attraction. The force field $\vec{f}_{ai}$ is the negative gradient of the potential [19] with a weight of $w_{i}$. When $\|\vec{r}_{ai}\|=\sigma_{a}$, $\vec{f}_{ai}=\textbf{0}$, so the formulation of interaction fields are:
@@
where $\vec{e}_{ai}$ is a unit vector from $R_{a}$ to $R_{i}$. The amplitude of $\|\vec{f}_{ai}\|$ is depicted in Fig. 2(c). With an increase in $w_{i}$, the impact of neighbors becomes more pronounced. The optimal value of $w_{i}$ can be obtained with in Section IV-B, and its robustness is demonstrated in simulations. Additionally, when $0<\|\vec{r}_{ai}\|<\sigma_{a}$, the interaction force behaves as a repulsion to avoid path interference. When $\sigma_{a}<\|\vec{r}_{ai}\|<3\sigma_{a}$, the force is gravitational for cooperative perception. When $\|\vec{r}_{ai}\|\geq 3\sigma_{a}$, the force decays almost to 0. We define $\|\vec{r}_{ai}\|\in(\sigma_{a},3\sigma_{a})$ as the pass band for observation fusion. Neighboring robots outside the pass band are either too close or too far away. In the application, it is essential to construct a segmentation function $\vec{F}_{ai}$ that performs differently under varying distances. The final expression for interaction forces is defines as:
@@
where $P_{ni}$ is the probability mapped from the information flow $\Sigma_{i}$. The significance of $P_{ni}$ is the probability to follow the neighbor in future paths. $P_{ni}$ will be high if the neighbor $R_{i}$ observes few obstacles towards the target of $R_{a}$. This will result in a strong attractive effect, thus increasing the likelihood of $R_{a}$ to follow it. In contrast, if the location of $R_{i}$ is obstructed by obstacles, $P_{ni}$ will be lower, leading to a divergence effect. In such cases, $R_{a}$ will be more likely to take a different path, which avoids future detours that would otherwise be caused by its observational limitation. Given that mobile robot navigation only involves left and right turns, $P_{ni}$ needs to be differentiated according to the relative orientation of the neighbors:
$P_{ni}=\begin{cases}\bm{\sigma}[\mathbf{1}^{\top}\cdot(\mathbf{x}_{ni}+\mathbf{w}_{n})]&\theta(\vec{r}_{ai})\geq 0\\ 1-\bm{\sigma}[\mathbf{1}^{\top}\cdot(\mathbf{x}_{ni}+\mathbf{w}_{n})]&\theta(\vec{r}_{ai})<0\end{cases},$ (8)
where $\mathbf{x}_{ni}$ and $\mathbf{w}_{n}$ are both 2D vectors, $\bm{\sigma}[\cdot]$ is the sigmoid function, and $\theta(\cdot)\in[-\pi,\pi]$ denotes the angle of a vector. The elements of $\mathbf{x}_{ni}$ are the estimates of passable distances in the navigation scope ($Q_{1}$ and $Q_{2}$) towards the target direction of $R_{a}$, and $\mathbf{w}_{n}$ are the corresponding distance thresholds to follow.
III-B2 Individual intelligence
The objective of $R_{a}$ is to reach its target in the shortest way. It evaluates potential planning directions based on its environmental observations. We use a force $\vec{F}_{a}$ to quantify its intrinsic motivation. The direction of the force is the same as the target direction $\vec{e}_{a}$. When an obstacle is detected in the navigation scope, $R_{a}$ should select a direction for circumnavigation. In this context, the amplitude of $\vec{F}_{a}$ is designed as the probability of searching its path on the left side:
$\vec{F}_{a}=P_{a}\vec{e}_{a},$ (9)
where $P_{a}$ is the probability determined by current observation:
$P_{a}=\begin{cases}\bm{\sigma}[\mathbf{W_{0}^{\top}}\mathbf{x_{0}}+b_{0}]&\text{obstacle exist}\\ 0.5&\text{other cases}\end{cases},$ (10)
with the estimated area of open spaces $\mathbf{x_{0}}$, and the weight vector $\mathbf{W_{0}}$ in the four quadrants. The bias entry $b_{0}$ is usually set to zero without a priori. The values of $\mathbf{W_{0}}$ assign weights to observations in both the navigation scope and the peripheral vision. In our experiments, they are equally weighted with odd symmetry left and right.
III-B3 Fusion optimization
The interaction fields provide additional information to identify more efficient paths with fewer detours. When $R_{a}$ encounters an obstacle, it fuses its own intelligence $\vec{F}_{a}$ and the resultant of $\vec{F}_{ai}$ for subsequent planning directions:
$P$ $=\frac{1}{Z}\left(\|\vec{F}_{a}\|+\sum\nolimits_{\vec{F}_{ai}\cdot\vec{e}_{ai}>0}\|\vec{F}_{ai}\|\right),$ (11)

### v3

@@ Abstract
This paper proposes a lightweight decentralized solution for multi-robot coordinated navigation with cooperative perception. First, we introduce a rapid way to process sensory data, thus obtaining safe directions and key environmental features. Then, an information flow is created to facilitate real-time perception sharing over wireless ad-hoc networks. Consequently, the environmental uncertainties of each robot are reduced by interaction fields that deliver complementary information. Finally, path optimization is achieved in a probabilistic way, enabling self-organized coordination with effective convergence, divergence, and collision avoidance. Our method is fully interpretable and ready for deployment without gaps. Comprehensive simulations and real-world experiments demonstrate reduced path redundancy, robust performance across various tasks, and minimal demands on computation and communication.
Note to Practitioners
Local perception is the information source for robot navigation in unknown environments. The extended perception at the swarm level provides complementary information for each robot. This can optimize robot paths and achieve coordinated navigation in a distributed way. However, this purpose is currently unachievable due to the high demands of communication and computation on board. This study reduces these two demands to the extreme by Fast Fourier Transform (FFT), digital filtering, probabilistic fusion, and swarm interaction forces. At the same time, our fully interpretable method avoids sim-to-real gaps and can be deployed directly on ground mobile robots. Simulation and real-world experiments outperform a state-of-the-art approach. Future work aims to extend our method to aerial robots as drones.
@@ I Introduction
In recent years, multi-robot systems have shown great potential in wild [1], industry [2], and indoor services [3]. Self-organized swarms stand out for their better automation and wider application. Despite significant progress, most solutions only use each robot’s local perception independently for navigation, while few studies have so far utilized real-time cooperative perception of the swarm to optimize the planning process. Consequently, robots’ coordination is often confined to leader-follower [4, 5] or collision avoidance [1, 6] due to limited information flow within swarm systems, in which robots often encounter unexpected obstructions or unnecessary detours.
The advantages of cooperative perception are obvious: a wider field of view, more details of the unknown environment, and that lead to better paths with fusion optimization. The challenges include but are not limited to informative environmental representation, decision-making with fused perception, on-board processing latency, and minimal unreliable communication. The first two are method requirements, and the latter two are hardware constraints.
@@
This paper proposes Nav-SCOPE: Swarm COoperative PErception and coordinated Navigation. This is an extremely lightweight systematic solution that addresses the above contradiction. In this system, each robot is used as a telescope of neighboring robots for an extended field of view. When robots decide their future paths, they will use the fused perception to choose better directions with fewer detours or conflicts. Experimental results show self-organized paths and better swarm automation. To the best of our knowledge, Nav-SCOPE is the first interpretable navigation solution with real-time cooperative perception, and it does not require a mapping process or external localization. The paper makes the following contributions:
@@
  • The navigation process is detailed for better swarm coordination and less path redundancy. Parameters can be determined directly before deployment.
  • Comprehensive simulations in challenging environments demonstrate that Nav-SCOPE outperforms a previously state-of-the-art approach. Real-world experiments show that our method can be applied robustly without gaps.
This article is organized as follows. Section II presents the related work. Section III gives the preliminaries of our work. In Section IV, the formulation of Nav-SCOPE is detailed, including sensor data processing, information flow from cooperative perception, coordinated navigation, and path optimization. The necessary steps for the deployment of Nav-SCOPE are introduced in Section V. Simulation and real word experiments are presented in Section VI and Section VII, respectively. Finally, Section VIII presents the conclusions of this study.
@@ II Related Works
Numerous studies have focused on local planning since the advent of mobile robots. Autonomous navigation was first realized in individual robots. Reactive strategies, such as Potential Fields (PFs) [7, 8] and Bugs [9, 10], are direct and lightweight solutions. Planners based on voxel maps [11, 12] can provide better local trajectories with more environmental details. Learning-based methods such as reinforcement learning [13] and guidance planning [14] have also excelled for their automation and become popular in recent years. These methods do not require a pre-built map, as a Simultaneous Localization and Mapping (SLAM) [15] process before planning execution is not suitable for many tasks and cases.
@@
As robot systems grow to fleets and swarms, critical issues of coordination between robots need to be solved. Centralized control is a typical way to regulate robot paths. For example, [8] and [16] used central computing stations to achieve flock flight on multiple quadrotors. The former is a typical solution based on optimization, while the latter is achieved by reinforcement learning. However, these approaches require ideal communication and external localization, making them vulnerable in scaled-up scenarios [17]. [3, 2] used received signal strength indication (RSSI) to achieve distributed navigation, but still needed external beacons to provide a priori environmental information, which limits the scalability of the algorithm.
@@
Decentralized systems can be more versatile in large unknown environments. The more information utilized, the better performance a system would have. Currently, it is common to use mutual positioning and trajectory sharing to achieve self-organization and swarm navigation. Consequently, recent studies have provided two types of solutions. (i) Leader-follower structures achieve synergic formation towards the same goal, such as leading robot following [18], formation control [4, 5] and target tracking [19, 20], which are mainly based on distance keeping and barycentric coordinates. (ii) Collision avoidance structures optimize robot paths to reduce negative interactions in various tasks, such as spatial-temporal optimization of trajectories [21, 1], and similar approaches performed by neural networks [22, 6]. A typical state-of-the-art (SOTA) solution is EGO-Swarm-v2 [1], which was tested in wild bamboo forests.
@@
Cooperative perception offers external environmental information to each robot, thereby holding significant potential to enhance swarm navigation and improve their performance. However, it is challenging to form an information flow among robots, as well as to take advantage of it. [23, 24] proposed Graph Neural Network (GNN) pipelines to share perceptual features, while the swarm planning process was not yet considered. [25] achieved a systematic GNN framework from perception to planning execution, but robot actions were limited to discrete on grid maps, thereby suffering from sim-to-real gaps. In contrast, Nav-SCOPE is designed to solve these challenges and can be immediately deployed in real-world applications.
III Preliminaries
We consider a swarm of ground mobile robots $R_{i},i\in\{1,2,\dots,N\}$ operating in unknown cluttered environments. Each robot has its local perception within the maximum range of $R$. The sensory data is defined as a digital depth sequence $s_{i}(n)$, with $M$ uniform samples within its field of view (FOV):
$s_{i}(n)=\mathbf{d}\left[\theta_{n}\in\Theta\right]\leq R,n=1,2,\dots,M,$ (1)
where $\Theta$ is the FOV, and $\mathbf{d}$ is the sequent depth array obtained from LiDARs or cameras, with each item corresponding to a distance in a certain direction $\theta_{n}$. Consider the physical dimension $r_{0}$ of a robot, a safety margin $r$ is used to keep the robot from obstacles, as shown in Fig. 2. The FOV is divided into four quadrants: navigation scope ($Q_{1}$, $Q_{2}$), and peripheral vision ($Q_{3}$, $Q_{4}$). They are used to evaluate the occupation of the future route. The scope angle $\alpha$ and planning distance $l_{th}$ are determined by geometric relationships:
@@ III Preliminaries
The necessary and sufficient condition of a safe direction can be expressed as a safety window in $Q_{1}$ and $Q_{2}$:
$\theta_{m}\text{ is safe}\iff\text{for }n\in(m-\frac{\alpha}{2},m+\frac{\alpha}{2}),s_{i}(n)\geq l_{th}.$ (3)
In this study, each robot can advance or rotate left and right towards a safe direction. The task of the swarm is to navigate safely and coordinately towards targets. The objective of Nav-SCOPE is to provide forward directions to robots, thus minimizing swarm path lengths with the least collision and resource consumption.
Figure 2: Safe margin and four observation quadrants. If there is no obstacle within the orange scope, the robot can keep a safety margin $r$ from the obstacle $B$.
IV Formulation of Nav-SCOPE
The system architecture of Nav-SCOPE is illustrated in Fig. 1. In this section, we will first introduce the FFT process and our filter design. Then, the information flow among swarm robots is formed. Finally, coordinated navigation facilitated by cooperative perception is elaborated.
Figure 3: Sensor data processing framework (right) and an illustrative example (left) in local coordinate. Blue boxes identify safe directions and green boxes extract environmental features. The feature excludes narrow spaces beneath the robot dimension, while maintains prospective directions for reconstruction. The computation complexity in (ii) and (iv) is $O(m\log m)$, while others are less than $O(m)$. The volume of environmental feature $\Sigma_{i}$ is usually less than $(2\Theta/\alpha)$.
IV-A FFT Process and Filter Design
In the perception stage, Nav-SCOPE scans to evaluate all directions of $s_{i}(n)$ within the FOV. The evaluation process can be performed rapidly by digital filtering based on Fast Fourier Transform (FFT), which was demonstrated in a previous study [26], since convolution in the time domain is equivariant to multiplication in the frequency domain. The overall complexity is $O(m\log m)$ in computation, and an illustration of the framework is shown in Fig. 3 (i)-(iv). The evaluation of this stage serves two functions: first, to identify safe routes and second, to extract key features of the surrounding environment.
The design of digital filters is critical for FFT filtering. In order to achieve the two functions discussed above, we propose two types of filters, respectively. First, to identify safe routes subject to robot dimensions $r_{0}$ and planning distance $l_{th}$, a normalized rectangular window can be designed to evaluate the occupancy of each direction, which is mathematically equivalent to (3). The size of the window $T_{c}$ corresponds to $\alpha$ in the physical world:
$h_{1}(n)=\frac{1}{T_{c}}R_{T_{c}}(n-\tau),n=1,2,\dots,M,$ (4)
where $T_{c}=M(\alpha/\Theta)$ refers to the window size normalized by the digital resolution, and $\tau=\frac{M-1}{2}$ is the group delay. This filter operates on short-term perception within the planning distance, so $s_{i}(n)$ is truncated to $l_{th}$ before being filtered. The safety condition (3) is further simplified with the filtered sequence $y_{i1}(n)$:
$\theta_{m}\text{ is safe}\iff y_{i1}(m)\geq l_{th},$ (5)
and we obtain the set of safe directions $\Theta_{s}=\{\theta_{m}|y_{i1}(m)\geq l_{th}\}$ for local planning, e.g., the sector in Fig. 3 (iii).
Second, Nav-SCOPE is required to evaluate environmental occupancy at the level of perceptual range in the long term. A low-pass filter can effectively extract overall spacial tendency, while filtering out minor details for subsequent compression. Consequently, another filter $h_{2}$ is designed to extract the key features of open space and potential obstructions according to the same window size $T_{c}$:
$h_{2}(n)=2f_{c}\text{sinc}\left(2f_{c}(n-\tau)\right)w_{M}(n,\beta),$ (6)
where $f_{c}=1/T_{c}$, $\text{sinc}(x)=sin(\pi x)/(\pi x)$, and $w_{M}(n,\beta)$ is a Kaiser window [27] with parameter $\beta$. As illustrated in Fig. 3(iv), the filtered sequence $y_{i2}(n)$ contains several sectors. The centers of these sectors are prospective pathways in long-term planning, and the areas are approximate estimates of open spaces.
IV-B Information Flow from Cooperative Perception
The information flow within the swarm is formed by distributed sharing of robot perception. In order to minimize communication loads, Nav-SCOPE performs informative encoding at transmitter, and rapid reconstruction at receiver. The centers and borders of the sectors in $y_{i2}(n)$ indicate possible open spaces and potential obstacles, respectively. Therefore, they are key perceptual features encoded for data transmission. The features can be obtained by calculating the local extrema of $y_{i2}(n)$, which are then masked by safe directions $\Theta_{s}$ to filter out narrow spaces relative to robot size (Fig. 3 (vi)-(vii)):
$\mathbf{\Sigma}_{i}=\left\{(\phi_{i,j},d_{i,j})|\Delta y_{i2}(j)\Delta y_{i2}(j+1)<0\cap\Theta_{s}\right\},$ (7)
where $\phi_{i,j}$ and $d_{i,j}$ are the directions and distances of the extrema. Consequently, we get perceptual features of robots only at the byte level. At the receiver, the directions $\phi_{i,j}$ should be calibrated according to the relative pose of two robots. Then, we use simple linear interpolation for approximate reconstruction, as shown in Fig. 3 (viii), and the information flow within the swarm is established.
In our applications in unknown, cluttered environments, robots are often obstructed by obstacles. The objective of the information flow is to deliver complementary observation of the surrounding environment. Consequently, an effective expansion should keep an overlap between the perceptual regions of two robots. This condition ensures correlation and accessibility between these observations. Hence, robots only connect to the neighboring robots within the FOV, which does not need a strongly connected network, but is time-varying and forms local groups spontaneously. In addition, the information flow of Nav-SCOPE is event-triggered and in an ad hoc way. When a robot needs cooperative perceptions from its neighbors, it will request data transmissions, and then the information flows in a one-way direction to the robot in need. In summary, the transmission mechanism is as follows:
@@ IV-B Information Flow from Cooperative Perception
Thus, the data are minimized while sufficient information is maintained, and the complexity of transmission scales linearly with the number of observed neighbors. The reduction in communication resource improves reliability under rigorous conditions or unreliable networks.
Figure 4: Detailed cases of coordination in simulation. (a) Convergence case. (b) Divergence case. (c) A detailed illustration of fusion optimization in (b). $R_{a}$ has similar probabilities in either directions before fusion. The perception of it’s neighbor $R_{b}$ in quadrant $Q_{1}$ provides strong probability of obstruction. Consequently, $R_{a}$ choose to navigate rightwards. (d) Distributed ultra-wide band (UWB) model of each robot for communication and mutual positioning.
IV-C Coordinated Navigation and Path Optimization
The objective of coordinated navigation is to optimize robot paths in a distributed way, combining local perception with swarm information flow. Here, neighboring robots perform as telescopes to provide complementary information in terms of the environment, in other words, help to find open spaces with fewer obstacles towards the target. We use forces to indicate the level of impact.
IV-C1 Individual intelligence
Robots use their local perception to evoke a tendency in a certain direction within $\Theta_{s}$. Considering a robot $R_{a}$, we use a force $\vec{F}_{a}$ to quantify its motivation. Note that there are two types of turning directions, namely left and right, for ground robots, the amplitude of $\vec{F}_{a}$ is designed as the probability of searching its path on the left side. If there is no obstruction within the planning distance $l_{th}$, $R_{a}$ will head towards the target direction $\vec{e}_{a}$ with the same probability on both sides. Otherwise, when an obstacle is detected in the navigation scope, $R_{a}$ should select a direction for circumnavigation:
$\vec{F}_{a}=P_{a}\vec{e}_{a},$ (8)
where $P_{a}$ uses a sigmoid function $\bm{\sigma}(\cdot)$ to map local perception into the probability domain:
$P_{a}=\begin{cases}\bm{\sigma}[\mathbf{W_{0}^{\top}}\mathbf{x_{0}}+b_{0}]&\exists\text{ obstacle}\\ 0.5&\text{other cases}\end{cases},$ (9)
with open space areas $\mathbf{x_{0}}$ from the reconstructed local $\Sigma_{a}$, and the weight vector $\mathbf{W_{0}}$ in the four quadrants. The bias entry $b_{0}$ is usually set to zero without a priori. The values of $\mathbf{W_{0}}$ assign weights to observations in both the navigation scope and the peripheral vision. In our experiments, they are equally weighted with odd symmetry left and right.
IV-C2 Interaction fields
They are connections among robots that deliver information flow, regulate swarm coordination, and achieve path optimization. Robots observe their neighbors within the FOV in an ad-hoc way, so we define the interaction field only exerted from local observed neighbors. In Nav-SCOPE, robot connectivity is time-varying according to the visibility and distances of neighboring robots. Consider a robot $R_{a}$ with its neighbor $R_{i}$. Consequently, the influence of neighbor $R_{i}$ is closely related to its observation features $\Sigma_{i}$ and their relative distances $\|\vec{r}_{ai}\|$. The potential $V_{ai}$ of the interaction combines the effect of coordination convergence $V^{att}_{ai}$ and collision avoidance $V^{rep}_{ai}$ [28]:
@@ IV-C2 Interaction fields
where $\sigma_{a}$ is the reference distance, marking the transition from repulsion to attraction. The force field $\vec{f}_{ai}$ is the negative gradient of the potential with a weight of $w_{i}$. When $\|\vec{r}_{ai}\|=\sigma_{a}$, $\vec{f}_{ai}=\textbf{0}$, so the formulation of interaction fields is as follows:
@@
where $\vec{e}_{ai}$ is a unit vector from $R_{a}$ to $R_{i}$. The amplitude of $\|\vec{f}_{ai}\|$ is depicted in Fig. 5(b). With an increase in $w_{i}$, the impact of neighbors becomes more pronounced. The optimal value of $w_{i}$ can be obtained in Section V-C, and its robustness is demonstrated in simulations. Furthermore, when $0<\|\vec{r}_{ai}\|\leq\sigma_{a}$, the interaction force behaves as a repulsion to avoid path interference. When $\sigma_{a}<\|\vec{r}_{ai}\|\leq 3\sigma_{a}$, the force is gravitational for cooperative perception. When $\|\vec{r}_{ai}\|>3\sigma_{a}$, the force decays almost to 0. We define $\|\vec{r}_{ai}\|\in(\sigma_{a},3\sigma_{a}]$ as the pass band for observation fusion. Neighboring robots outside the pass band are either too close or too far away. A segmentation function $\vec{F}_{ai}$ is designed to perform differently at varying distances:
@@
where $P_{ni}$ is the probability of following the neighbor in future paths. $P_{ni}$ will be high if the neighbor $R_{i}$ observes few obstacles towards the target of $R_{a}$. This will result in a strong attractive effect, thus increasing the likelihood of $R_{a}$ to follow it. In contrast, if the location of $R_{i}$ is blocked by obstacles, $P_{ni}$ will be lower, leading to a divergence effect. In such cases, $R_{a}$ will be more likely to take a different path, which avoids future detours that would otherwise be caused by its observational limitation. The definition of $P_{ni}$ is differentiated by the quadrant in which the neighbor is observed, in accordance with left searching in (9):
$P_{ni}=\begin{cases}\bm{\sigma}[\mathbf{1}^{\top}\cdot(\mathbf{x}_{ni}+\mathbf{w}_{n})]&R_{i}\in\{Q_{1},Q_{3}\}\\ 1-\bm{\sigma}[\mathbf{1}^{\top}\cdot(\mathbf{x}_{ni}+\mathbf{w}_{n})]&R_{i}\in\{Q_{2},Q_{4}\}\end{cases},$ (14)
where $\mathbf{x}_{ni}$ indicate the estimates of the passable distances towards the target, and $\mathbf{w}_{n}$ are the distance thresholds of following. They are both 2D vectors corresponding to the scope quadrants ($Q_{1}$ and $Q_{2}$), after receiving and calibrating the neighboring perception $\Sigma_{i}$.
IV-C3 Fusion optimization
The interaction fields provide additional information to identify more efficient paths with fewer detours. When $R_{a}$ encounters an obstacle, it combines its own intelligence $\vec{F}_{a}$ and the resultant of $\vec{F}_{ai}$ for subsequent planning directions:
$P_{dir}$ $=\frac{1}{Z}\left(\|\vec{F}_{a}\|+\sum\nolimits_{\vec{F}_{ai}\cdot\vec{e}_{ai}>0}\|\vec{F}_{ai}\|\right),$ (15)